# Boosting Visual-Language Models by Exploiting Hard Pairs

### v2

@@ 3.3 HNML: hard negative margin loss
Figure 2: Hard Negative Margin Loss (HNML). Hard negative pairs are closer to the positive than the normal negative pairs.
Figure 2: Hard Negative Margin Loss (HNML). Hard negative pairs are closer to the positive than the normal negative pairs.
Figure 3: Further training CLIP with Hard Pairs. For text-image pairs within a batch, we sample corresponding hard data from the preprocess hard pair set.
@@
$\ell_{margin}=\frac{1}{|B|}\sum_{j\in B}\max\big{(}0,sim(I_{i},T_{j})-\min_{j^{\prime}\in\mathcal{H}^{p}_{i}}\{sim(I_{i},T_{j^{\prime}})\}\big{)},$ (6)
@@
$\ell_{finetune}=\ell_{CLIP}+\gamma\ell_{margin}.$ (7)
@@ 4 Experiments
In the experiments, we conduct a comprehensive empirical investigation to evaluate the efficacy of Helip in improving zero-shot classification, image-text retrieval, and linear probing performances for existing vision-language models, in Section 4.2. In Sections 4.3 and 4.4, we investigate Helip’s performance with scaled training data as well as its robustness over noisy datasets. We provide detailed empirical studies on Hard Positive Mining (HPM) and Hard Negative Mining with Margin Loss (HNML) in Sections 4.6 and 4.7.
@@ 4.1 Experimental setup
Training datasets. We used open-source datasets, including Conceptual Captions 3M (CC3M) (Sharma et al., 2018) and Conceptual Captions 12M (CC12M) (Changpinyo et al., 2021), and two 15M subsets of the YFCC100M dataset: v1, collected by Radford et al. (2021), and v2, collected by  Li et al. (2022b). The combined datasets of CC3M, CC12M, and YFCC15M v1, which we denote it as Open29M following the term used in prior work (Li et al., 2022b), were not completely obtained due to expired urls. In addition, we independently sampled 7.5M and 8M subsets from the noisier data source, LAION-5B (Schuhmann et al., 2022), labeled as LAION7.5M and LAION8M. These datasets, while smaller than the 400 million pair dataset used in CLIP’s original study (Radford et al., 2021), are well-suited for the data and computational resources we have. Furthermore, they have been widely used in benchmark evaluations for various studies on language-image pretraining, as noted in works by  Goel et al. (2022); Li et al. (2022b) and Mu et al. (2022).
Downstream datasets. We primarily evaluate the effectiveness of Helip using zero-shot image classification, linear probing, and zero-shot image-text retrieval. In addition to commonly used ImageNet  (Deng et al., 2009), CIFAR10, and CIFAR100 (Krizhevsky et al., 2009), we also verify the performance on 7 fine-grained classification datasets including Caltech101 (Fei-Fei et al., 2004), Food101 (Bossard et al., 2014), Sun397 (Xiao et al., 2010), Flowers102 (Nilsback & Zisserman, 2008), CUB (Wah et al., 2011), Stanford Cars (Krause et al., 2013) and FGVC Aircraft Maji et al. (2013). The zero-shot image-text retrieval task uses MS-COCO (Lin et al., 2014) and Flickr30K (Plummer et al., 2015).
Implementation details. Our experiments are conducted across three distinct architectures: ResNet-50, ViT-B/16, and ViT-B/32, tailored to various datasets and pretrained models. Specifically, for loading the pretrained CLIP model on CC3M and CC12M, the ResNet-50 is used as the image encoder. Besides, to align with existing checkpoints established by  Mu et al. (2022), we use ViT-B/16 for SLIP model experiments on CC3M and CC12M, respectively. And, we use ViT-B/32 for pretraining on YFCC15M v1, v2, and Open29M datasets to ensure fair comparison with the results reported in  Li et al. (2022b). Furthermore, for the SLIP and DECLIP models, we adapt the pretrained parameters from the publicly available resources [^1] https://github.com/facebookresearch/SLIP, https://github.com/Sense-GVT/DeCLIP. The input resolution of the image encoder is 224 $\times$ 224 and the maximum context length of the text encoder is 77. All of our experiments are conducted on 8 V100 GPUs with a batch size of 128 for ViT-B/16 models, and a batch size of 512 for ResNet-50 models and ViT-B/32 models. The dimension of the image and text embeddings is 1024 for ResNet-50 models and 512 for ViT-B/16 and ViT-B/32 models. We set $\tau=0.5$, $\gamma=1$ and $p=1$ for all the experiments by default. Automatic mixed-precision is used to save GPU memory. To keep the model from overfitting, we use early stopping if there is no performance gain in 5 epochs. Note, to reflect that our method is designed to work with few assumptions on encoder, we used encoders pretrained over a single-modal source rather than multimodally pretrained ones when preparing hard negative pairs. Specifically, we used an unsupervised pre-trained vision transformer, DINO VITs8 (Caron et al., 2021), and a Sentence Transformer (SentenceT) (Reimers & Gurevych, 2019) to encode text. For DINO VITs8, the embedding size is 384, while for SentenceT, it is 768.
@@ 4.2 Main results and discussion
Zero-shot classification. We compare the zero-shot performances of the CLIP, SLIP, DECLIP, and those models finetuned by Helip on CC3M, CC12M, YFCC15M and Open29M. We denote the models finetuned by Helip as CLIP-Helip , SLIP-Helip , and DECLIP-Helip respectively. Table 1 demonstrates that models fine-tuned by Helip consistently outperform their counterparts. Specifically, for models pretrained on the CC3M dataset, Helip boosts the ImageNet zero-shot classification accuracy of the CLIP model from 19.04% to 19.86%. Additionally, on the SLIP model, a performance improvement of over 13% is observed, achieving an accuracy of 26.05%. We additionally include two baseline methods: CYCLIP (Goel et al., 2022) and CLOOB (Fürst et al., 2021) for reference. For CC12M pretraining, we used the SLIP checkpoints released by Mu et al. (2022). On ImageNet, SLIP-Helip has a 4.47% higher zero-shot accuracy than its counterpart. Due to the lack of openly accessible parameters for DECLIP on the CC3M and CC12M datasets, our analysis focused on comparing DECLIP with DECLIP-Helip over the YFCC15M v2 dataset. In this context, we present the performance of the SLIP and DECLIP models, as pretrained and released by  Li et al. (2022b). The result was obtained by using their evaluation pipeline, denoted with *{*}*. Note, the templates are important for zero-shot tasks. Consequently, for a fair comparison, our analysis and conclusions primarily rely on results obtained from our evaluation pipeline, which is same as the approach used by OpenCLIP. Due to space constraints, we provide more information about the baselines in Appendix B. Notably, both SLIP and DECLIP showed improvements with Helip, averaging increases of 15.49% and 6.74%, respectively. Further, to demonstrate Helip’s sustained efficacy across larger datasets, we assessed CLIP and CLIP-Helip on Open29M. The original CLIP model, upon training with the Open29M dataset, achieves its bset performance at the 18th epoch, achieving a zero-shot accuracy of 42.32% on ImageNet. Extending the training duration by two epochs resulted in a marginal decline in accuracy to 42.25%. Notably, our Helip method instantly boost the performance of the existing CLIP (checkpoint saved at 18th epoch) from 42.32% to 46.33% with just one additional training epoch.
@@
${}^{\dagger}$ (Radford et al., 2021) 19.04 33.06 13.77
${}^{\dagger}$-Helip 19.86 34.05 14.13
@@
CC12M ${}^{\dagger}$ (Radford et al., 2021) 30.27 51.07 21.94
${}^{\dagger}$-Helip 32.05 52.27 24.51
@@
29M ${}^{\dagger}$ (Radford et al., 2021) 42.32 71.98 42.73
${}^{\dagger}$-Helip 46.33 77.97 48.33
@@
Zero-shot fine-grained classification. By leveraging hard image-text pairs in contrastive learning, Helip amplifies the discriminative capability of the CLIP model’s visual embedding. This improvement proves valuable in classification, particularly for fine-grained datasets. Our evaluation on 7 fine-grained classification datasets (Table 2) reveals that SLIP-Helip boosts the zero-shot accuracy of CC3M and CC12M pretrained SLIP on Caltech101 by 12.88% and 3.95% respectively. Both CLIP and SLIP models witness consistent improvements with their Helip counterparts.
@@
Linear probing. The linear probing task trains a randomly initialized linear classifier on the feature extracted from the frozen image encoder on the downstream dataset. To accomplish this, we train the logistic regression classifier using scikit-learn’s L-BFGS implementation (Pedregosa et al., 2011), with maximum 1,000 iterations on those 7 datasets. For each dataset, we search for the best regularization strength factor on the validation set over 45 logarithmically spaced steps within the range 1e-6 to 1e+5. Experimental results in Table 3 demonstrate that both CLIP-Helip and SLIP-Helip have consistent improvements over their counterparts on almost all 7 datasets. Note that on CC12M, SLIP-Helip performs marginally better on 5 out of 7 datasets. It’s probably because the self-supervision of SLIP (Mu et al., 2022) within the visual modal can be beneficial for learning fine-grained visual embedding, while SLIP-Helip doesn’t include image self-supervision during the training. In addition, we did not match the training batch size as SLIP (Mu et al., 2022) because of resource limitations. A combination of Helip and image self-supervision and larger training batch size may be a potential direction for achieving better linear probe performance.
@@
Pretraining Dataset Method COCO Flickr30K
R@1 $\uparrow$ R@5 $\uparrow$ R@1 $\uparrow$ R@5 $\uparrow$
CC3M CLIP 14.4 34.1 31.7 56.0
CLIP-Helip 17.8 39.8 35.4 61.0
SLIP 22.3 45.6 39.6 68.6
SLIP-Helip 23.4 48.3 41.8 69.6
CC12M CLIP 26.9 52.6 47.2 74.3
CLIP-Helip 27.8 54.3 48.2 75.4
SLIP 39.0 66.0 65.4 90.1
SLIP-Helip 39.4 67.2 66.2 89.7
Table 4: Zero-shot image-text retrieval results on MSCOCO and Flickr. $\uparrow$ indicates higher is better. Combining with Helip, CLIP and SLIP show better performance.
4.3 Performance of HELIP with Scaled Training Data
@@
4.4 Performance of HELIP on noisy dataset

### v3

@@ 3.3 HNML: Hard Negative Margin Loss
Figure 2: Hard Negative Margin Loss (HNML). Hard negative pairs (i.e., the golden retriever) are closer to the positive than the normal negative pairs.
@@
$\ell_{\text{margin}}=\frac{1}{|B|}\sum_{j\in B}\max\big{(}0,sim(I_{i},T_{j})-\min_{j^{\prime}\in\mathcal{H}^{p}_{i}}\{sim(I_{i},T_{j^{\prime}})\}\big{)},$ (6)
@@
$\ell_{\text{finetune}}=\ell_{\text{CLIP}}+\gamma\ell_{\text{margin}}.$ (7)
@@
Figure 3: Continuous training CLIP with Hard Pairs. For text-image pairs within a batch, we sample corresponding hard data from the preprocess hard pair set.
@@ 4 Experiments
In Section 4.2, we empirically investigate Helip’s efficacy in improving zero-shot classification, image-text retrieval, and linear probing performance of existing vision-language models. Section 4.3 examines Helip’s performance with scaled training data. We provide in-depth studies on Hard Positive Mining (HPM) and Hard Negative Mining with Margin Loss (HNML) in Sections 4.4 and 4.5, respectively. Discussions on robustness over noisy datasets and additional empirical analyses of hard mining methods are deferred to the appendix.
@@ 4.1 Experimental Setup
Training Datasets. We used open-source datasets including CC3M (Sharma et al., 2018), CC12M (Changpinyo et al., 2021), and two 15M subsets of YFCC100M: v1 (Radford et al., 2021) and v2 (Li et al., 2022b). The combined datasets—CC3M, CC12M, and YFCC15M v1 (Open29M (Li et al., 2022b))—were not fully obtained due to expired URLs. Additionally, we sampled 7.5M and 8M subsets from the noisier LAION-5B (Schuhmann et al., 2022), labeled LAION7.5M and LAION8M. Although smaller than the 400M pairs used in CLIP’s original study (Radford et al., 2021), these datasets suit our computational resources and have been widely used in language-image pretraining benchmarks (Goel et al., 2022; Li et al., 2022b; Mu et al., 2022).
Downstream Datasets. We evaluate Helip using zero-shot image classification, linear probing, and zero-shot image-text retrieval. Beyond ImageNet (Deng et al., 2009), CIFAR10, and CIFAR100 (Krizhevsky et al., 2009), we assess performance on seven fine-grained classification datasets: Caltech101 (Fei-Fei et al., 2004), Food101 (Bossard et al., 2014), Sun397 (Xiao et al., 2010), Flowers102 (Nilsback and Zisserman, 2008), CUB (Wah et al., 2011), Stanford Cars (Krause et al., 2013), and FGVC Aircraft Maji et al. (2013). For zero-shot image-text retrieval, we use MS-COCO (Lin et al., 2014) and Flickr30K (Plummer et al., 2015). Implementation details are provided in the appendix.
@@ 4.2 Main Results and Discussion
Zero-Shot Classification. We evaluate the zero-shot performance of CLIP, SLIP, and DECLIP models, along with their Helip-fine-tuned versions (CLIP-Helip, SLIP-Helip, and DECLIP-Helip), on the CC3M, CC12M, YFCC15M, and Open29M datasets. Table 1 shows that models with Helip consistently outperform their counterparts. Specifically, on the CC3M dataset, Helip boosts the ImageNet zero-shot accuracy of CLIP from 19.04% to 19.86% and improves SLIP by over 13% to 26.05%. We include two baseline methods, CYCLIP (Goel et al., 2022) and CLOOB (Fürst et al., 2021), for reference. Using SLIP checkpoints from Mu et al. (2022) pretrained on CC12M, SLIP-Helip achieves a 4.47% higher zero-shot accuracy on ImageNet than SLIP. Since DECLIP parameters for CC3M and CC12M are unavailable, we compare DECLIP and DECLIP-Helip on the YFCC15M v2 dataset, also presenting SLIP and DECLIP models pretrained by Li et al. (2022b) evaluated with their pipeline (denoted with {*}). Because templates significantly impact zero-shot tasks, we use our evaluation pipeline (the same as OpenCLIP) for fair comparison; further baseline details are in Appendix B. Both SLIP and DECLIP show improvements with Helip, averaging increases of 15.49% and 6.74%, respectively. To demonstrate Helip’s efficacy on larger datasets, we evaluated CLIP and CLIP-Helip on Open29M. The original CLIP model reaches its best zero-shot ImageNet accuracy of 42.32% at epoch 18. Applying Helip boosts this to 46.33% with just one additional epoch, whereas continued training with the original CLIP loss slightly decreases accuracy to 42.25%.
@@
CLIP (Radford et al., 2021) 19.04 33.06 13.77
CLIP-Helip 19.86 34.05 14.13
@@
CC12M CLIP (Radford et al., 2021) 30.27 51.07 21.94
CLIP-Helip 32.05 52.27 24.51
@@
29M CLIP (Radford et al., 2021) 42.32 71.98 42.73
CLIP Cont. Train 42.25 71.72 42.66
CLIP-Helip 46.33 77.97 48.33
@@
Zero-Shot Fine-Grained Classification. Utilizing hard image-text pairs in contrastive learning, Helip enhances the discriminative power of CLIP’s visual embeddings, benefiting fine-grained classification tasks. As shown in Table 2, SLIP-Helip improves zero-shot accuracy on Caltech101 by 12.88% and 3.95% for models pre-trained on CC3M and CC12M, respectively. Both CLIP and SLIP models consistently improve when augmented with Helip. The above results indicate that the embedding space becomes tighter when using hard pairs in contrastive loss.
@@
Pretraining Dataset Method COCO Flickr30K
R@1 $\uparrow$ R@5 $\uparrow$ R@1 $\uparrow$ R@5 $\uparrow$
CC3M CLIP 14.4 34.1 31.7 56.0
CLIP-Helip 17.8 39.8 35.4 61.0
SLIP 22.3 45.6 39.6 68.6
SLIP-Helip 23.4 48.3 41.8 69.6
CC12M CLIP 26.9 52.6 47.2 74.3
CLIP-Helip 27.8 54.3 48.2 75.4
SLIP 39.0 66.0 65.4 90.1
SLIP-Helip 39.4 67.2 66.2 89.7
Table 3: Zero-shot image-text retrieval results on MSCOCO and Flickr. $\uparrow$ indicates higher is better.
Linear Probing. The linear probing task trains a randomly initialized linear classifier on the feature extracted from the frozen image encoder on the downstream dataset. We train the logistic regression classifier using scikit-learn’s L-BFGS implementation (Pedregosa et al., 2011), with maximum 1,000 iterations on those 7 datasets. For each dataset, we search for the best regularization strength factor on the validation set over 45 logarithmically spaced steps within the range 1e-6 to 1e+5. Experimental results in Table 4 demonstrate that both CLIP-Helip and SLIP-Helip have consistent improvements over their counterparts on almost all 7 datasets. Note that on CC12M, SLIP-Helip performs marginally better on 5 out of 7 datasets. It’s probably because the self-supervision of SLIP (Mu et al., 2022) within the visual modal can be beneficial for learning fine-grained visual embedding, while SLIP-Helip doesn’t include image self-supervision during the training. And we did not match the training batch size as SLIP (Mu et al., 2022) because of resource limitations. A combination of Helip and image self-supervision with larger batch size may be a potential direction for achieving better linear probe performance.
@@
4.3 HELIP with Scaled Training Data
@@
Figure 5: Hard pairs from HPM and fastHPM. FastHPM produces high-quality hard pairs that compete with HPM.
4.4 Impact of Hard Negative Margin Loss
We examine the impact of hard negative margin loss (HNML) on the SLIP model’s performance. Specifically, we analyze the SLIP model pre-trained on CC3M when further trained with both HPM and HNML, compared to training without HNML. We evaluated zero-shot classification performance across datasets including ImageNet, CIFAR-100, CIFAR-10, Caltech-101, Food-101, and SUN397. As detailed in Table 5, the SLIP model augmented with HPM and HNML outperforms the baseline SLIP and SLIP + HPM models by 4.51 and 3.27 points, respectively. Notably, the model achieves better performance on CIFAR-10 without HNML. We hypothesize that HNML enhances the discriminative power of learned representations by incorporating class distance as a cost metric. Therefore, for classification tasks with a larger number of subclasses, employing HNML during training can improve performance.
ImageNet CF10 CF100 Caltech101 Food101 Sun397 Avg.
SLIP 23.00 65.61 34.69 54.01 16.03 29.20 37.09
wo HNML 24.94 69.44 36.35 64.07 16.51 30.91 40.37
w HNML 26.05 68.18 37.77 66.89 17.05 33.68 41.60
Table 5: SLIP continous training with and without hard negative margin loss.
4.5 Delving into Hard Pair Mining
Performance Comparison between HPM and FastHPM. A comparison was made between the zero-shot performances of SLIP models, further trained with hard pairs obtained from both HPM and fastHPM. This comparison, conducted under three different settings, was summarized in Table 6. Additionally, we established subsets $\widetilde{\mathcal{D}}_{i}$ of sizes 3M and 6M, and accordingly denoted Helip with these subset sizes as Helip-3M and Helip-6M. Table 6 shows that the zero-shot performances of Helip-3M and Helip-6M remain competitive with the global HPM hard pair mining approach. These findings suggest that fastHPM offers an efficient hard pair mining strategy without compromising performance and has the potential to scale up in larger pre-training datasets.
Imagenet CIFAR10 CIFAR100
SLIP 41.17 81.30 53.68
Helip- 3M 45.07 82.42 55.22
Helip- 6M 44.98 81.64 56.62
Helip- Full 45.64 82.31 53.79
Table 6: Zero-shot performance of SLIP with HELIP on CC12M with hard pairs from HPM and fastHPM.
Visual insights into HPM and FastHPM. We took the initiative to visualize the hard pairs as identified by the aforementioned three methods. Within Figure 5, the leftmost image-text pairing is earmarked as the target. The pairs in the primary row are those selected via HPM. The subsequent rows, specifically the second and third, present image-text pairings identified by the 6M fastHPM and the 3M fastHPM methods, respectively. Through a comparative visualization, it’s evident that the hard pairs pinpointed by fastHPM bear a significant resemblance to the target pair. For readers keen on delving deeper, we’ve provided an extended set of visualization outcomes in Appendix J.
Computational Time Analysis. Table 7 provides a comparison of the computational time required by HPM and fastHPM. The hard negative pairs preparation times listed were measured on 8 V100 GPUs, with the exception of the * symbol, which was measured on a single V100 GPU. Given its efficiency and the performance similarities observed in Table 6, fastHPM emerges as a compelling alternative to the full HPM method.
CC3M CC12M YFCC15M
Helip- 3M - 2h18min 3h27min
Helip- 6M - 5h3min 6h19min
Helip- Full 1h9min 9h11min 17h41min
5 Conclusion
In this work, we present Helip, a framework that enhances pre-trained CLIP models by more effectively utilizing their original training datasets. Helipoffers a cost-effective and easily integrable solution for improving existing models without extensive retraining or additional data. Specifically, it treats each text-image pair as a point in the joint vision-language space and identifies hard pairs, those that are close together, using the Hard Pair Mining (HPM) strategy. Furthermore, to efficiently leverage these hard pairs, we introduce the Hard Negative Margin Loss (HNML). Empirically, we found that Helipboosts the performance of existing checkpoints within a few epochs of continuous training. Evaluations across various benchmarks, including zero-shot classification, image-text retrieval, and linear probing, demonstrate the effectiveness and efficiency of our method. These findings highlight that in the era of large-scale models and datasets, performance improvement can be achieved not only by collecting more data or scaling up models, but also by intelligently maximizing the utility of the data we already have.
6 Limitations
While Helip demonstrates significant improvements over existing CLIP models, several limitations should be acknowledged.
Dependence on Dataset Quality and Diversity.
The effectiveness of Helip is inherently tied to the quality and diversity of the original training dataset. If the dataset contains limited variability or is biased toward certain domains, the benefits of mining hard pairs may diminish. In scenarios where datasets are very small, the selection of meaningful challenging pairs becomes more difficult, potentially limiting the overall performance gains.
Reusing Original Training Data.
Helip leverages the original pretraining data without incorporating additional external data sources. While this approach reduces the need for costly data collection, it might also inherits biases present in the original dataset.
Limited Evaluation on Diverse Datasets.
The empirical evaluations of Helip are conducted on specific datasets like CC3M, CC12M, YFCC15M and subsets of LAION. While these are widely accepted benchmarks, the generalizability of its improvements to real-world applications requires further evaluation.
7 Acknowledgments
This material is based upon work supported by the Air Force Office of Scientific Research under award number FA2386-24-1-4011. This research is also partially supported by the Singapore Ministry of Education Academic Research Fund Tier 1 (Award No. T1 251RES2207).
Additional support for Minbin Huang and Hong Cheng is provided by project #MMT-p2-23 of the Shun Hing Institute of Advanced Engineering at The Chinese University of Hong Kong, and by the Research Grants Council of the Hong Kong Special Administrative Region, China (No. CUHK 14217622).
The authors gratefully acknowledge all of these sources of support. Any opinions, findings, and conclusions or recommendations expressed in this publication are those of the authors and do not necessarily reflect the views of the funding agencies.
Appendix A Appendix: Algorithm
Input: Hard pairs number per sample $k$
Pretrained unimodal vision model: $f_{\text{text}}$
Pretrained unimodal vision model: $f_{\text{image}}$
Dataset $\mathcal{D}=\{(x_{1}^{I},x_{1}^{T}),(x_{2}^{I},x_{2}^{T}),\cdots,(x_{N}^{I},x_{N}^{T})\}$
Threshold for visual and textual modality $\tau_{I}$ and $\tau_{T}$
for $i\in[1,N]$ do
       $\mathbf{s}\leftarrow[0,0,\cdots,0]^{\top}\in\mathbb{R}^{N}$
       $I_{i}\leftarrow f_{\text{image}}(x_{i}^{I})$
       $T_{i}\leftarrow f_{\text{text}}(x_{i}^{T})$
       for $j\in[1,N]$ do
             $I_{j}\leftarrow f_{\text{image}}(x_{j}^{I})$
             $T_{j}\leftarrow f_{\text{text}}(x_{j}^{T})$
             $\vec{S}^{I}_{j}\leftarrow\frac{I_{i}\cdot I_{j}}{\left\|I_{i}\right\|_{2}\cdot\left\|I_{j}\right\|_{2}}\,\textbf{if}\,\frac{I_{i}\cdot I_{j}}{\left\|I_{i}\right\|_{2}\cdot\left\|I_{j}\right\|_{2}}>\tau_{I}\,\textbf{else}\,0$
             $\vec{S}^{T}_{j}\leftarrow\frac{T_{i}\cdot T_{j}}{\left\|I_{i}\right\|_{2}\cdot\left\|T_{j}\right\|_{2}}\,\textbf{if}\,\frac{T_{i}\cdot T_{j}}{\left\|T_{i}\right\|_{2}\cdot\left\|T_{j}\right\|_{2}}>\tau_{T}\,\textbf{else}\,0$
             $\mathbf{s}_{j}\leftarrow\vec{S}^{I}_{j}\cdot\vec{S}^{T}_{j}$
       end for
      $\mathcal{H}_{i}\leftarrow\operatorname*{arg\,max}(\mathbf{s},k)$
       if $\exists j\in\mathcal{H}_{i}$, $\mathbf{s}_{j}=0$ then
             $\mathcal{H}_{i}=\emptyset$  #  Indicate noise sample
end for
Algorithm 1 Hard Pair Mining (HPM)
Input: Hard pairs number per sample $k$
Pretrained unimodal vision model: $f_{\text{text}}$
Pretrained unimodal vision model: $f_{\text{image}}$
Dataset $\mathcal{D}=\{(x_{1}^{I},x_{1}^{T}),(x_{2}^{I},x_{2}^{T}),\cdots,(x_{N}^{I},x_{N}^{T})\}$
Threshold for visual and textual modality $\tau_{I}$ and $\tau_{T}$
Candidate pool size $C$
for $i\in[1,N]$ do
       Uniformly $C$ samples from Dataset $\mathcal{D}$, $\overline{\mathcal{D}}_{i}=\{(x_{1}^{I},x_{1}^{T}),(x_{2}^{I},x_{2}^{T}),\cdots,(x_{C}^{I},x_{C}^{T})\}$
       $\mathbf{s}\leftarrow[0,0,\cdots,0]^{\top}\in\mathbb{R}^{N}$
       $I_{i}\leftarrow f_{\text{image}}(x_{i}^{I})$
       $T_{i}\leftarrow f_{\text{text}}(x_{i}^{T})$
       for $j\in[1,C]$ do
             $I_{j}\leftarrow f_{\text{image}}(x_{j}^{I})$
             $T_{j}\leftarrow f_{\text{text}}(x_{j}^{T})$
             $\vec{S}^{I}_{j}\leftarrow\frac{I_{i}\cdot I_{j}}{\left\|I_{i}\right\|_{2}\cdot\left\|I_{j}\right\|_{2}}\,\textbf{if}\,\frac{I_{i}\cdot I_{j}}{\left\|I_{i}\right\|_{2}\cdot\left\|I_{j}\right\|_{2}}>\tau_{I}\,\textbf{else}\,0$
             $\vec{S}^{T}_{j}\leftarrow\frac{T_{i}\cdot T_{j}}{\left\|I_{i}\right\|_{2}\cdot\left\|T_{j}\right\|_{2}}\,\textbf{if}\,\frac{T_{i}\cdot T_{j}}{\left\|T_{i}\right\|_{2}\cdot\left\|T_{j}\right\|_{2}}>\tau_{T}\,\textbf{else}\,0$
             $\mathbf{s}_{j}\leftarrow\vec{S}^{I}_{j}\cdot\vec{S}^{T}_{j}$
       end for
      $\mathcal{H}_{i}\leftarrow\operatorname*{arg\,max}(\mathbf{s},k)$
       if $\exists j\in\mathcal{H}_{i}$, $\mathbf{s}_{j}=0$ then
             $\mathcal{H}_{i}=\emptyset$  #  Indicate noise sample
end for
Algorithm 2 fast Hard Pair Mining (fastHPM)
Input: $\mathcal{D}=\{(x_{1}^{I},x_{1}^{T}),(x_{1}^{I},x_{1}^{T}),\cdots,(x_{N}^{I},x_{N}^{T})\}$
Hard Pair Mining algorithm, $\text{HPM}()$  # or the $\text{fastHPM}()$
Pretrained unimodal vision model: $f_{\text{text}}$
Pretrained unimodal vision model: $f_{\text{image}}$
Pretrained contrastive language-image model $\{\phi_{\text{image}},\phi_{\text{text}}\}$
hyperparameters:
 Hard pairs number $k$
 Hard negative margin strength $\gamma$
 Sampled hard negatives number $p$
 Learning ratio $\eta$
 Batch size $b$
 Training iteration number $E$
 Visual and textual modality threshold $\tau_{I}$ and $\tau_{T}$
Output: CLIP model $\{\phi_{\text{image}},\phi_{\text{text}}\}$
$\mathcal{H}\leftarrow\text{HPM}(\mathcal{D},f_{\text{text}},f_{\text{image}},k,\tau_{I},\tau_{T})$
for $iter\in[1,E]$ do
       $B\leftarrow\{z_{1},\ldots,z_{b}\}\stackrel{{\scriptstyle\text{ i.i.d. }}}{{\sim}}Uniform(\mathcal{D})$
       for $z_{i}\in B$ do
             $\mathcal{H}_{i}^{p}\leftarrow\{z_{i},\ldots,z_{p}\}\stackrel{{\scriptstyle\text{ i.i.d. }}}{{\sim}}Uniform(\mathcal{H}_{i})$
             $\overline{B}\leftarrow B\cup\mathcal{H}_{i}^{p}$
       end for
      Compute $\text{Compute loss}~{}\ell_{\text{finetune}}\text{, Equation (6), with samples}~{}\overline{B}$ $\phi_{\text{image}}\leftarrow\phi_{\text{image}}+\eta\cdot\partial_{\phi_{\text{image}}}\ell_{\text{finetune}}$ $\phi_{\text{text}}\leftarrow\phi_{\text{text}}+\eta\cdot\partial_{\phi_{\text{text}}}\ell_{\text{finetune}}$
end for
Appendix B Appendix: Discussion about baselines
Appendix C Appendix: Analysis of the Impact of Subset Size on Hard Pair Selection in FastHPM
Figure 6: The average selection criteria values for hard pairs mined by FastHPM with different subset sizes.
In the comparison of HPM and FastHPM detailed in Section 4.5, we explore the efficacy of using 3M and 6M subset sizes of the CC12M dataset in FastHPM for mining hard pairs. The result, Table 6, shows that with a reduced subset size as small as 3 million entries, mining hard pairs and further training with these pairs can boost CLIP to achieve competitive performance with full set for mining.
In this section, we delve deeper into the analysis of hard pairs mined by FastHPM across varying subset sizes. Based on the selection criteria defined by FastHPM (Equation 5), we denote the selection criteria value as $\widetilde{S}^{I}(x^{I}_{i},\mathcal{H}^{\star}_{i}(j))^{\top}\widetilde{S}^{T}(x^{T}_{i},\mathcal{H}^{\star}_{i}(j))$. Here, $\mathcal{H}^{\star}_{i}(\cdot)$ represents a pair within the set of hard pairs $\mathcal{H}^{\star}_{i}$, mined by FastHPM for a specified target pair $i$ under a given subset size. Additionally, the $j$ in $\mathcal{H}^{\star}_{i}(j)$ indicates the $j$-th hard pair within the set $\mathcal{H}^{\star}_{i}$. Note, a higher selection criteria value signifies a harder mined pair.
We present the average selection criteria values for top-k hard pairs in Figure 6. As depicted by the grey horizontal line, the average selection criteria values for the top-20 hard pairs selected by FastHPM-1.5M, the top-40 by FastHPM-3M, and the top-80 by FastHPM-6M all approximate 0.477. This figure indicates that a further reduction in the subset size might necessitate adjustments to the number of hard pairs sampled to preserve quality. For instance, in our experiments detailed in Table 6, we uniformly sampled hard pairs for training from the top 50 for Helip-3M. As Figure 6 suggests, a sampling range of 10 for Helip-1M might be effective. Particularly, considering that Helipsignificantly boosted the pre-trained models with just an additional training epoch, as discussed in Section 4.2, selecting one hard pair for each target pair from a pool of 10 will be feasible.
Appendix D Appendix: Implementation Details
Our experiments are conducted across three distinct architectures: ResNet-50, ViT-B/16, and ViT-B/32, tailored to various datasets and pretrained models. Specifically, for loading the pretrained CLIP model on CC3M and CC12M, the ResNet-50 is used as the image encoder. Besides, to align with existing checkpoints established by  Mu et al. (2022), we use ViT-B/16 for SLIP model experiments on CC3M and CC12M, respectively. And, we use ViT-B/32 for pretraining on YFCC15M v1, v2, and Open29M datasets to ensure fair comparison with the previous results Li et al. (2022b). Furthermore, for the SLIP and DECLIP models, we adapt the pretrained parameters from the publicly available resources [^2] https://github.com/facebookresearch/SLIP, https://github.com/Sense-GVT/DeCLIP. The input resolution of the image encoder is 224 $\times$ 224 and the maximum context length of the text encoder is 77. All of our experiments are conducted on 8 V100 GPUs with a batch size of 128 for ViT-B/16 models, and a batch size of 512 for ResNet-50 models and ViT-B/32 models. The dimension of the image and text embeddings is 1024 for ResNet-50 models and 512 for ViT-B/16 and ViT-B/32 models. We set $\tau=0.5$, $\gamma=1$, $k=50$ and $p=1$ for all the experiments by default. Automatic mixed-precision is used to save GPU memory. To keep the model from overfitting, we use early stopping if there is no performance gain on ImageNet zero-shot accuracy in 5 epochs. It is worth noting that using zero-shot classification performance on ImageNet as a criterion for early stopping is a commonly used practice for the training of CLIP (Radford et al., 2021; Mu et al., 2022).
To reflect that our method is designed to work with few assumptions on encoder, we used encoders pretrained over a single-modal source rather than multimodally pretrained ones when preparing hard negative pairs. Specifically, we used an unsupervised pre-trained vision transformer, DINO VITs8 (Caron et al., 2021), and a Sentence Transformer (SentenceT) (Reimers and Gurevych, 2019) to encode text. For DINO VITs8, the embedding size is 384, while for SentenceT, it is 768.
Appendix E Appendix: Performance of HELIP on noisy dataset